OpenStudy (anonymous):

Group Theory. Isomorphic Groups on the Set R. G is the set of the real numbers with the operation x*y=x+y+1. Find an isomorphism f: R->G and show that it is an isomorphism.

OpenStudy (kinggeorge):

Do you know what the identity of G is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Well, maybe....-(x+y+1)

OpenStudy (kinggeorge):

In fact, the identity, would be -1. You can solve for it by looking at the equation \(xy=x+y+1=y\). Subtract \(y\), and you get \(x+1=0\implies x=-1\). Since any isomorphism maps identity to identity, you know your map sends \(0\mapsto-1\). This should suggest an isomorphism. In particular,\[\begin{aligned}f:R&\longrightarrow G\\x&\longmapsto x-1\end{aligned}\]

OpenStudy (kinggeorge):

I think this map should work. You just need to prove it's an isomorphism.

OpenStudy (anonymous):

Thank you for your help! i've struggled with this problem all weekend.

OpenStudy (kinggeorge):

No problem.